# Spatial Association Between Near-Misses and Accident Blackspots in Sydney, Australia: A Getis-Ord $G_{i}^{*}$ Analysis

### v1

@@ VI Conclusion
This study investigated the spatial relationship between reported road crashes and a high-severity near-miss proxy (High-G events) in Sydney using Bivariate Local Moran’s I (LISA) over a 400m grid. The analysis confirmed significant spatial clustering of crashes and revealed distinct local patterns of association between the two indicators. Concordant hotspot areas were identified where high crash rates coincided with statistically high neighbouring High-G rates (HH pattern, $825$ cells). However, no areas showed a statistically significant pattern where both indicators were concurrently low (LL pattern, 00 significant cells found at $p<0.05$). Crucially, discordant areas were also found: locations with high crashes but statistically low neighbouring High-G events (HL pattern, $91$ cells), and a substantial number of locations ($2681$ cells) with low crashes despite statistically high neighbouring High-G events (LH pattern).
These findings demonstrate that while High-G events have spatial distribution does not perfectly mirror that of reported crashes. Characterizing the different pattern areas (HH, HL, LH) using Point of Interest (POI) data revealed significant differences in land use context (e.g., HH areas showing different POI profiles compared to HL or LH areas, potentially highlighting varying environmental contributors to risk). The discordant patterns are particularly valuable: HL areas ($91$) may point to crash causes not captured well by the High-G metric, while the numerous LH areas ($2681$) suggest locations with frequent risky events where crashes are currently mitigated or under-reported, warranting proactive investigation and monitoring. Limitations include data accuracy, the specific definition and proxy nature of High-G events, and potential Modifiable Areal Unit Problem (MAUP) effects from the grid analysis. Future work should incorporate dynamic traffic flow and detailed infrastructure data to better explain these complex spatial patterns and advance proactive road safety strategies.
Future research should focus on incorporating dynamic traffic variables, detailed infrastructure data, and potentially driver behavior information to build more comprehensive models explaining the observed spatial patterns of concordance and discordance. Network-based analyses and qualitative case studies of specific HL and LH locations could provide further findings into the factors mitigating or exacerbating risk. Ultimately, a multi-faceted approach combining leading and lagging indicators with rich contextual data is essential for advancing proactive road safety management.

### v2

@@ VI Conclusion
This study investigated the spatial relationship between reported road crashes and a high-severity near-miss proxy (High-G events) in Sydney using LISA over a 400m grid. The analysis confirmed significant spatial clustering of crashes and revealed distinct local patterns of association between the two indicators. Concordant hotspot areas were identified where high crash rates coincided with statistically high neighbouring High-G rates (HH pattern, $825$ cells). However, no areas showed a statistically significant pattern where both indicators were concurrently low (LL pattern, 0 significant cells found at $p<0.05$). Crucially, discordant areas were also found: locations with high crashes but statistically low neighbouring High-G events (HL pattern, $91$ cells), and a substantial number of locations ($2681$ cells) with low crashes despite statistically high neighbouring High-G events (LH pattern).
These findings demonstrate that the spatial distribution of High-G events shows statistically significant concordance (p¡0.05) with historical crash patterns in $825$ cells (HH clusters), while identifying $2681$ additional high-risk areas (LH clusters) where frequent near-misses occur without corresponding crashes. This indicates High-G data identifies additional locations requiring monitoring beyond those identified solely by historical crash data, potentially revealing areas where frequent risky maneuvers occur without corresponding crash history.
Characterizing the different pattern areas using POI data revealed significant differences in land use context between these spatial patterns. The discordant patterns are particularly valuable: HL areas ($91$) may point to crash causes not captured well by the High-G metric, while the numerous LH areas ($2681$) suggest locations with frequent risky events where crashes are currently mitigated or under-reported, warranting proactive investigation and monitoring. Limitations include data accuracy, the specific definition and proxy nature of High-G events. Future work should incorporate dynamic traffic flow and detailed infrastructure data to better explain these complex spatial patterns and advance proactive road safety strategies.
Future research should focus on incorporating dynamic traffic variables, detailed road infrastructure data, and potentially localized driver behaviour information to build more comprehensive models explaining the observed spatial patterns of concordance and discordance. Network-based analysis and qualitative case studies of specific HL and LH locations could provide further findings into the factors mitigating or exacerbating risk. Ultimately, an approach combining leading and lagging indicators with contextual data is essential for advancing proactive road safety management.